# The NHIM development scenario in an asymmetric binary system of dwarf galaxies

### v1

@@ 4.2 The horizontal horseshoe
The properties of homoclinic and heteroclinic connections can be illustrated best by the construction of horseshoes in some appropriate Poincaré map. We have already seen in the previous subsection that horizontal heteroclinic trajectories help a lot to illustrate heteroclinic connections between the various NHIMs. In addition, we can only plot Poincaré maps restricted to some invariant 2-dof subsystem. Therefore, we construct now a Poincaré map for the horizontal subsystem and illustrate the creation of heteroclinic connections between the NHIMs in this 2-dimensional map. For the construction of this Poincaré map, we use the intersection condition $y=0$ with negative orientation. Remember that it is the same intersection condition which we have used before in the construction of Fig.2. Therefore, we can relate important structures in the horizontal Poincaré map with particular periodic orbits contained in Fig.2 by just comparing the $x$ coordinates. The domain of the horizontal Poincaré map is the $x-p_{x}$ plane.
@@
As already mentioned before, the only horizontal trajectories within the NHIMs over the collinear Lagrange points are the respective horizontal Lyapunov orbits. And in the horizontal Poincaré map, they are represented by fixed points on the line $p_{x}=0$. Their $x$ coordinate can be read off Fig.2. We use the energy value -0.043 because this is the highest energy where the three horizontal Lyapunov orbits lh1, lh2, and lh3 still exist, are still normally hyperbolic and still act as boundary curves of the NHIMs, as explained in section 3. Therefore, it is clear that the corresponding fixed points in the horizontal Poincaré map are hyperbolic fixed points. In the Poincaré plot, these three fundamental fixed points are marked by the black open circles and are called lh1, lh2 and lh3 respectively, exactly equal to the Lyapunov orbits which they represent.
@@
We calculate initial segments of the stable and unstable manifolds $W^{s}(lhj)$ and $W^{u}(lhj)$ of the three fundamental fixed points and plot them into the $x-p_{x}$ plane, i.e. into the domain of the horizontal map. The result is Fig.11. From the trajectory plots in the horizontal position space shown in the previous subsection, it is obvious that we have frequent violations of the transversality of the intersection condition $y=0$. Therefore, the stable and unstable manifolds are torn into many disjoint fragments. In the plot $W^{s}(lh1)$, $W^{u}(lh1)$, $W^{s}(lh2)$, $W^{u}(lh2)$, $W^{s}(lh3)$ and $W^{u}(lh3)$ are plotted in the colours orange, green, red, blue, magenta and cyan, respectively. The black curves are the boundary of the energetically accessible region of the $x-p_{x}$ plane at E=0.043E=-0.043.
@@
Figure 11: Horseshoe at E=0.043E=-0.043 in the invariant horizontal subsystem built up by the stable and unstable manifolds of the horizontal Lyapunov orbits sitting over the collinear Lagrange points. In the plot $W^{s}(lh1)$, $W^{u}(lh1)$, $W^{s}(lh2)$, $W^{u}(lh2)$, $W^{s}(lh3)$, and $W^{u}(lh3)$ are plotted in the colours orange, green, red, blue, magenta, and cyan, respectively. The fixed points representing the Lyapunov orbits are marked by black open circles. The black curves are the boundary of the energetically accessible region at E=0.043E=-0.043. Some KAM islands are included as grey curves.
@@
In the figure, we make an interesting observation. Between the region around the Lagrange points $L_{1}$ and $L_{2}$, there exists an amazingly large KAM island which reaches almost to the energetic boundary of the domain of the map. This makes heteroclinic connections to the NHIM2 relatively difficult. In comparison, the KAM island between the regions around the Lagrange points $L_{1}$ and $L_{3}$ is a lot smaller and leaves more room for heteroclinic connections to NHIM3. This observation makes it understandable that NHIM1 has more coordination of its development scenario with NHIM3 than with NHIM2. Of course, for very high iterations of the initial segments of the stable and unstable manifolds, the ones from any one of the NHIMs accumulate against the ones from all other stable or unstable manifolds respectively of all invariant subsets of the dynamics. However, the corresponding heteroclinic trajectories might need a very long time to come from the direct neighbourhood of the initial invariant subset to the neighbourhood of the final invariant subset. On their way, these trajectories perform a lot of complicated intermediate loops.
It seems that for any efficient coordination of the development scenarios of NHIMs, we need heteroclinic trajectories which proceed along simple and fast trajectories between the direct neighbourhoods of these NHIMs.
@@ 4.3 Some remarks on transients and heteroclinic connections
The existence of the heteroclinic trajectories starting or ending on transient parts of decaying NHIMs might lead us to the following interesting speculation: The beginning of the decay of NHIM3 and the formation of the corresponding transient chaotic regions is synchronised with the formation of a large-scale chaotic region on NHIM1. Such a synchronisation is supported by the existence of heteroclinic trajectories like the one presented in Fig.9(b).
@@
After having seen numerical evidence for transient effects and their influence on the general dynamics, let us finish this section with some further remarks on the decay of NHIMs and the related appearance of transient behaviour in the region of the former NHIM. As long as the NHIM is a smooth 3-dimensional invariant surface of the flow in the 5-dimensional energy shell (or a corresponding 2-dimensional surface in the 4-dimensional domain of the map), its stable manifold is a smooth 4-dimensional surface in the energy shell (or a corresponding 3-dimensional surface in the map). It must have codimension 1 in the energy shell and also in the domain of the map. Then a 1-dimensional curve $C$ in an appropriate region of the energy shell close to the NHIM intersects the local branch of the stable manifold transversally. The curve $C$ can be chosen to lie in the domain of the map, and then it also serves in the domain of the map for the same purpose. This is the basic idea behind the search strategy for the NHIM as explained in all details in [20].
@@
Now imagine that the NHIM decays and that in some region only a lower-dimensional fractal collection of remnants of the NHIM remains. These remnants certainly contain unstable periodic orbits. The stable manifolds of these remnants have a lower dimension, i.e. higher codimension than the stable manifold of the former complete NHIM. Then, in general, the curve $C$ will no longer intersect exactly the stable manifolds of the remnants. Accordingly, the search strategy for the NHIM mentioned above will no longer find these lower-dimensional remnants. However, if the curve $C$ is chosen in an appropriate region, it still contains points leading to a high time delay in the neighbourhood of the remnants. This is the explanation of the transient effects. To find again the stable manifolds of such lower-dimensional remnants of NHIMs, it would be necessary to scan initial conditions not only along a 1-dimensional curve but on a higher-dimensional surface, where the dimension of this surface of initial conditions should be larger than or at least equal to the codimension of the stable manifold we want to catch.
5 Indicator functions
@@
The phase space indicator functions are useful tools to analyse and get insights into the dynamics of systems with many degrees of freedom. A review of the topic and several examples can be found in [30, 31, 32]. In the present study, we are going to calculate the indicator based on the delay time. This quantity is an appropriate indicator function to study the phase space of open scattering systems [20, 24]. Some recent studies of the phase space and the transport using indicator functions for a similar model in 2 dimensions are in the reference [33].
@@
To calculate the delay time indicator, we select the initial conditions on the canonical conjugate plane $xp_{x}$ with $y=0$, $z=0$, and $p_{z}=0$ for different values of the energy $E$, see the Fig.12. The white regions on the plots correspond to forbidden regions for the value of $E$. The lines with maximal values of the delay time indicator correspond to intersections of the initial conditions with the stable and unstable manifolds of the NHIMs, the yellow smooth regions are the intersections with the KAM stable tori, and the green regions correspond to initial conditions that escape fast to infinity. As the value of the $E$ increases, the structure of the homoclinic tangle changes. Let us notice that the initial conditions for the panel Fig.12(a) and Fig.11 are the same. We appreciate the match between the stable and unstable manifolds of the NHIMs and the KAM stable islands on both types of plots using different visualisation techniques.
a) E=0.043E=-0.043
b) E=0.042E=-0.042
c) E=0.041E=-0.041
d) E=0.04E=-0.04
e) E=0.039E=-0.039
f) E=0.037E=-0.037
Figure 12: Delay time for different values of $E$ and $z=0$
To continue with the analysis of the phase space of the system, we consider initial conditions on the canonical conjugate plane $xp_{x}$, $y=0$, $p_{z}=0$ and a fixed value of the energy E=0.0043E=-0.0043, but different values of initial $z$. The Fig.13 shows the intersections of the heteroclinic tangles of the NHIMs and the KAM tori with the set of initial conditions. The phase space has a similar structure for different values of $z=1,2,3$. The structure of the set of heteroclinic intersections and KAM islands looks very similar to the case with $z=0$ in Fig.12(a). However, we can see a change in the forbidden region for $z=4$; the set of initial conditions has two components.
@@
Figure 13: Delay time indicator function for different values of $z$ and E=0.043E=-0.043
In the Fig.14, there is a fate map for the same initial conditions as Fig.12(a). For this plot, we only calculate the trajectories forward in time. The blue regions correspond to trajectories that escape to the interaction region with $x>0$, the green region to trajectories that stay trapped in the interaction region and the yellow region to trajectories that escape with $x<0$. We can see agreement betwee the delay time plot and the Poincaré map in Figs.12(a) and 11.
Figure 14: Fate map for E=0.043E=-0.043, $z=0$. The trajectories with yellow initial conditions escape to the left side, the dark blue ones escape to the right side, and the green ones stay trapped in the central region.
@@ 6 Conclusions and final remarks
The system analysed in the present article is an instructive example of Hamiltonian dynamics with 3 not symmetry-related NHIMs, and for the effects of coordination of the development scenario of the NHIMs. A new feature in the present article is the inclusion of tangentially transient behaviour on decaying NHIMs into the formation of the global scenario. We saw indications that also the decaying transient parts of NHIMs give important contributions to the heteroclinic tangle, i.e. to the global chaotic invariant set created by the NHIMs and their stable and unstable manifolds.
@@
In the study of the development scenario of the NHIMs and the comparison with the corresponding left-right symmetric system investigated in [21], we made the following interesting observation: Also, when we destroy the symmetry between the two outer NHIMs by the transition to different masses of the two dwarf galaxies, there remains a partial qualitative symmetry in the development scenarios of the two outer NHIMs. This partial symmetry includes the sequence of the important bifurcations and their order for the Lyapunov orbits contained in the individual NHIMs. But it does not include the coordination of the creation of chaos. Here, in this example, only NHIM1 and NHIM3 make a coordinated development into the direction of chaos; NHIM2 does not participate in this coordination. This happens even though we have heteroclinic trajectories between any pair of NHIMs. The horizontal Poincaré map presented in subsection 5.2 gives at least some indication of the relatively weak heteroclinic coupling of NHIM2 to the other two fundamental NHIMs of the system.
@@
It seems that NHIM1 and NHIM3 become susceptible to tangential instabilities at a similar energy interval, while NHIM2 is not yet ready for the development of large-scale chaos. Of course, in the symmetric case studied in [21], NHIM2 has exactly the same development scenario as NHIM3, and the two NHIMs develop large-scale chaos in exactly the same energy interval. The horizontal horseshoe plot of Fig.11 suggests that the large-scale KAM island over the potential minimum $P_{2}$ prevents a rapid and efficient heteroclinic connection between NHIM1 and NHIM2 along horizontal trajectories. In our collection of heteroclinic orbits, we do not have a single one which starts or ends on NHIM2, runs over the potential hole at $P_{2}$ and does not go to $z$ values far away from $z=0$. As we have seen in subsection 5.1, long and complicated trajectories can do it. As Figs.8(e) and 9(a) show, also trajectories going to the outside and circling around $L_{4}$ or $L_{5}$ can do it. But these long trajectories seem to be little efficient in the creation of coordination effects in the development scenarios between NHIMs. Of course, also trajectories using the vertical degree of freedom can establish heteroclinic connections, remember the trajectory from Fig.9(d). However, we have seen here (and also in other model systems) that in systems with a $z$ reflection symmetry and accordingly an invariant horizontal subsystem, the energy exchange between the vertical degree of freedom and the horizontal degrees of freedom is rather slow. This prevents an efficient coupling between the NHIMs via trajectories having a large part of their energy in the vertical motion. All these features prevent any efficient coupling of the development scenario of NHIM2 to the development scenario of the other two NHIMs.
@@
One important aspect of the development scenario, which we can not yet explain precisely, but which certainly deserves a further more fundamental investigation, is the following consideration: For stable and unstable manifolds of NHIMs, there holds a foliation theorem which loosely speaking says that these manifolds are the Cartesian product of a line with the internal NHIM structure. Thereby, these manifolds transport the internal NHIM structure along them to all regions of the phase space heavily influenced by the homoclinic/heteroclinic tangle of the NHIMs. As a consequence, the whole homoclinic/heteroclinic tangle is at least qualitatively similar to a Cartesian product of a lower-dimensional homoclinic/heteroclinic tangle (in our case, the horizontal horseshoe presented in subsection 5.2) and the internal NHIM structures. This property has been noticed before and investigated in more detail in [34]. Also, the system investigated in this reference is 3-dof and has, in a natural form, an invariant horizontal 2-dof subsystem. Such a tendency of higher-dimensional homoclinic/heteroclinic tangles to show an approximate product structure has also been observed by another group in [35, 36, 37]. This tendency seems to be common and needs a more profound explanation.
@@
The appearance of the tangential transient effects shows that, also after the beginning of its decay, the remnants of a NHIM have an effect on their environment in the phase space and on the formation of the homoclinic/heteroclinic tangle very similar to the effect the complete NHIM had before the start of the decay. The fragments and transient parts continue without heavy qualitative differences.
@@
The NHIMs sitting over the Lagrange points of the effective potential direct the transport of test particles over these saddle points. In the interpretation of our model as describing two dwarf galaxies in interaction, this model describes the exchange of stars between the two galaxies (motion over the central Lagrange point $L_{1}$) or the loss of stars from the binary galaxy system (motion over the outer Lagrange points $L_{2}$ and $L_{3}$). The projection of the unstable manifolds of the NHIMs into the position space then gives the position of tidal arms and tidal bridges going out from the dwarf galaxies. Of course, our model system could also be interpreted as a simple model for two nearby globular clusters in interaction. Here, it is interesting to mention that in the astronomical literature, there is a claim of observational indications for a tidal bridge between the interacting globular clusters NGC5024 and NGC5053 [38].

### v2

@@ 4.2 The horizontal horseshoe
The properties of homoclinic and heteroclinic connections can be illustrated best by the construction of horseshoes in some appropriate Poincaré map. We have already seen in the previous subsection that horizontal heteroclinic trajectories help a lot to illustrate heteroclinic connections between the various NHIMs. In addition, we can only plot Poincaré maps restricted to some invariant 2-dof subsystem. Therefore, we construct now a Poincaré map for the horizontal subsystem and illustrate the creation of heteroclinic connections between the NHIMs in this 2-dimensional map. For the construction of this Poincaré map, we use the intersection condition $y=0$ with negative orientation. Remember that it is the same intersection condition which we have used before in the construction of Fig.2. Therefore, we can relate important structures in the horizontal Poincaré map with particular periodic orbits contained in Fig.2 by just comparing the $x$ coordinates. The domain of the horizontal Poincaré map is the $xp_{x}$ plane.
@@
As already mentioned before, the only horizontal trajectories within the NHIMs associated with the collinear Lagrange points are the respective horizontal Lyapunov orbits. And in the horizontal Poincaré map, they are represented by fixed points on the line $p_{x}=0$. Their $x$ coordinate can be read off Fig.2. We use the Jacobi constant value $E_{J}=-0.043$ because this is the highest Jacobi constant where the three horizontal Lyapunov orbits $lh_{1}$, $lh_{2}$, and $lh_{3}$ still exist, are still normally hyperbolic and still act as boundary curves of the NHIMs, as explained in section 3. Therefore, it is clear that the corresponding fixed points in the horizontal Poincaré map are hyperbolic fixed points. In the Poincaré plot, these three fundamental fixed points are marked by the black open circles and are called $lh_{1}$, $lh_{2}$ and $lh_{3}$ respectively, exactly equal to the Lyapunov orbits which they represent.
@@
We calculate initial segments of the stable and unstable manifolds $W^{s}(lh_{i})$ and $W^{u}(lh_{i})$ of the three fundamental fixed points and plot them into the $xp_{x}$ plane, i.e. into the domain of the horizontal map. The result is Fig.11. From the trajectory plots in the horizontal position space shown in the previous subsection, it is obvious that we have frequent violations of the transversality of the intersection condition $y=0$. Therefore, the stable and unstable manifolds are torn into many disjoint fragments. In the plot $W^{s}(lh_{1})$, $W^{u}(lh_{1})$, $W^{s}(lh_{2})$, $W^{u}(lh_{2})$, $W^{s}(lh_{3})$ and $W^{u}(lh_{3})$ are plotted in the colours orange, green, red, blue, magenta and cyan, respectively. The black curves are the boundary of the energetically accessible region of the $xp_{x}$ plane at $E_{J}=-0.043$.
@@
Figure 11: Horseshoe at $E_{J}=-0.043$ in the invariant horizontal subsystem built up by the stable and unstable manifolds of the horizontal Lyapunov orbits associated with the collinear Lagrange points. In the Poincaré map, $W^{s}(lh_{1})$, $W^{u}(lh_{1})$, $W^{s}(lh_{2})$, $W^{u}(lh_{2})$, $W^{s}(lh_{3})$, and $W^{u}(lh_{3})$ are plotted in the colours orange, green, red, blue, magenta, and cyan, respectively. The fixed points representing the Lyapunov orbits are marked by black open circles. The black curves are the boundary of the energetically accessible region at $E_{J}=-0.043$. Some KAM islands around the minima of the effective potential energy $V_{eff}$ are included as grey curves.
@@
In the figure, we make an interesting observation. Between the region around the Lagrange points $L_{1}$ and $L_{2}$, there exists a large KAM island which reaches almost to the energetic boundary of the domain of the map. This makes heteroclinic connections to the NHIM $\mathcal{M}^{2}_{E_{J}}$ relatively difficult. In comparison, the KAM island between the regions around the Lagrange points $L_{1}$ and $L_{3}$ is a lot smaller and leaves more room for heteroclinic connections to NHIM $\mathcal{M}^{3}_{E_{J}}$. This observation makes it understandable that NHIM $\mathcal{M}^{1}_{E_{J}}$ has more coordination of its bifurcation scenario with NHIM $\mathcal{M}^{3}_{E_{J}}$ than with NHIM $\mathcal{M}^{2}_{E_{J}}$. Of course, for very high iterations of the initial segments of the stable and unstable manifolds, the ones from any one of the NHIMs accumulate against the ones from all other stable or unstable manifolds, respectively, of all invariant subsets of the dynamics. However, the corresponding heteroclinic trajectories might need a very long time to come from the direct neighbourhood of the initial invariant subset to the neighbourhood of the final invariant subset. On their way, these trajectories perform many complicated intermediate loops.
@@ 4.3 Some remarks on transients and heteroclinic connections
After having seen numerical evidence for transient effects and their influence on the general dynamics, let us finish this section with some further remarks on the break of NHIMs and the related appearance of transient behaviour in the region of the former NHIM. As long as the NHIM is a smooth 3-dimensional invariant surface of the flow in the 5-dimensional Jacobi constant manifold (or a corresponding 2-dimensional surface in the 4-dimensional domain of the map), its stable manifold is a smooth 4-dimensional surface in the Jacobi constant manifold (or a corresponding 3-dimensional surface in the map). It must have codimension 1 in the Jacobi constant manifold and also in the domain of the map. Then a 1-dimensional curve $C$ in an appropriate region of the Jacobi constant manifold close to the NHIM intersects the local branch of the stable manifold transversally. The curve $C$ can be chosen to lie in the domain of the map, and then it also serves in the domain of the map for the same purpose. This is the basic idea behind the search strategy for the NHIM as explained in all details in [22].
@@
Now imagine that, when we change the perturbation parameter, the NHIM loses its normal hyperbolicity in some region and breaks; only a lower-dimensional fractal collection of remnants of the NHIM remains. These remnants certainly contain unstable periodic orbits. The stable manifolds of these remnants have a lower dimension, i.e. higher codimension than the stable manifold of the former complete NHIM. Then, in general, the curve $C$ will no longer intersect exactly the stable manifolds of the remnants. Accordingly, the search strategy for the NHIM mentioned above will no longer find these lower-dimensional remnants. However, if the curve $C$ is chosen in an appropriate region, it still contains points leading to a high time delay in the neighbourhood of the remnants. This is the explanation of the transient effects. To find the stable manifolds of such lower-dimensional remnants of NHIMs again, it would be necessary to scan initial conditions not only along a 1-dimensional curve but on a higher-dimensional surface, where the dimension of this surface of initial conditions should be larger than or at least equal to the codimension of the stable manifold we want to catch.
5 Visualisation of the phase space using the technique of delay time
@@
To complement our study of the phase space of the system as the perturbation parameter grows we are going to use a phase space structure indicator based on the delay time to visualise the invariant objects in the phase space. The phase space indicator functions are a useful tool to analyse and get insights into the dynamics of systems with many degrees of freedom. A review of the topic and several examples can be found in the references [32, 33, 34]. Also, recent studies of the phase space using Lagrangian descriptors for a similar galactic model in 2-dof are presented in [35]. The delay time is one of the best indicators to study the phase space of open systems due to its simple interpretation and fast convergence; more information about it can be found in [22, 26].
Let us describe the construction of this phase space structure indicator function. First, we select a finite region $U$ of the 3-dimensional position space, which contains the position space projections of the important unstable invariant sets which we want to investigate. In our case, we think of the three important NHIMs. Let $X$ be any point in the phase space whose projection into the position space lies within $U$. Now we calculate the trajectory belonging to the initial point $X$ into the future direction of time and also into the past, in both time directions until the trajectory leaves the region $U$ for the first time. Let us say into the future direction, it happens for the positive time $t_{+}$, and for the past direction, it happens for the negative time $t_{-}$. Accordingly, the trajectory stays within the region $U$ for the time $\Delta t=t_{+}-t_{-}$. Then we define the time delay $t_{d}$ in $U$ as $t_{d}=\Delta t-t_{fm}$, where $t_{fm}$ represents uninteresting parts of the time in the asymptotic region where the particle behaves like a free particle. Of course, for practical calculations, we introduce a time limit $t_{l}$ and stop the calculation of the trajectory, if it does not escape from $U$ for a calculation time $t_{+}$ smaller than $t_{l}$ in future direction or a calculation time $t_{-}$ larger than $t_{l}$ in past direction.
To construct plots of the delay time indicator for our present system, we first chose as initial conditions the canonical plane $xp_{x}$ with $y=0$, $z=0$, and $p_{z}=0$, while the value of $p_{y}$ is determined by the fixed value of the Jacobi constant $E_{J}$, where the negative branch of $p_{y}$ is chosen in order to have the same orientation in $y$ direction as in Fig.2 and Fig.11. We select a fine grid of initial points on the $xp_{x}$ plane and calculate $t_{d}$ for each point. The result is presented in the Fig.12 for 6 different values of $E_{J}$ from the most interesting interval of $E_{J}$. The white regions on the plots correspond to energetically forbidden initial conditions for the given value of $E_{J}$. Yellow points indicate large values of the delay time $t_{d}$. These are points in stable KAM islands and also points lying on stable or unstable manifolds of unstable invariant sets, in particular, the ones of the important NHIMs. Blue colour indicates points which go to the asymptotic region rapidly in the future and also in the past.
@@
As the value of $E_{J}$ increases, the structure of the KAM tori and of the homoclinic and heteroclinic tangles changes but the NHIMs $\mathcal{M}^{1}_{E_{J}}$, $\mathcal{M}^{2}_{E_{J}}$, and $\mathcal{M}^{3}_{E_{J}}$ persist. Let us emphasise that the initial conditions for the plots on the panel Fig.12(a) and Fig.11 are the same. Therefore, we can appreciate the match between the stable and unstable manifolds of the NHIMs and the stable KAM islands on both types of plots, i.e. on Poincaré maps and on delay time indicator functions.
Comparing the panels in Fig.12 as the value of $E_{J}$ increases, we notice that the value of the delay time $t_{d}$ decrees on the region around the point $(x,p_{x})=(-5,0)$, close to the NHIMs $\mathcal{M}^{3}_{E_{J}}$. This fact is evidence that the NHIMs $\mathcal{M}^{3}_{E_{J}}$ is losing hyperbolicity and the transient chaotic region close to it is growing in size as the value of $E_{J}$ increases. This behaviour is expected when the value of $E_{J}$ increases far from the energy of the Lagrange point $L_{3}$, as we saw in previous sections. Similar behaviour of the delay time function indicator has seen in the references [26, 22].
a) $E_{J}=-0.043$
b) $E_{J}=-0.042$
c) $E_{J}=-0.041$
d) $E_{J}=-0.04$
e) $E_{J}=-0.039$
f) $E_{J}=-0.037$
Figure 12: Delay time indicator function in colour scale with initial conditions in the canonical plane $xp_{x}$, $y=0$, $z=0$, and $p_{z}=0$ for different values of $E_{J}$.
To continue with the analysis of the phase space of the system, we also consider initial conditions on the canonical plane $xp_{x}$, $y=0$, $p_{z}=0$ and again a fixed value of the Jacobi constant $E_{J}=-0.043$, but now for different initial values of $z$, namely $z=1,2,3,4$. For $z\neq 0$, the plane of initial conditions does not belong to a lower-dimensional invariant subsystem. Therefore, in general, trajectories do not return to this plane of initial conditions; the particle can move in all directions in the energy shell of the phase space.
The Fig.13 displays the tangles of the NHIMs and the KAM tori for these sets of initial conditions. The structure of the phase space looks qualitatively similar to the case with $z=0$ in Fig.12(a), we can recognise the intersections of the stable and unstable manifolds of the three NHIMs $W^{s/u}(\mathcal{M}^{1}_{E_{J}})$, $W^{s/u}(\mathcal{M}^{2}_{E_{J}})$, and $W^{s/u}(\mathcal{M}^{3}_{E_{J}})$ and the KAM tori. In addition, we can see a qualitative change of the energetically allowed region for $z=4$; here, the set of energetically accessible initial conditions has two disjoint components.
From Fig.13, we notice that the KAM tori associated with the point $P_{2}$ intersects the planes with different value of $z$, forming a barrier in the phase space. This KAM tori gives little room for the trajectories to travel between the neighbourhood around the NHIM $\mathcal{M}^{1}_{E_{J}}$ and the other two NHIMs $\mathcal{M}^{1}_{E_{J}}$ and $\mathcal{M}^{1}_{E_{J}}$, as we mention before in previous sections.
@@
Figure 13: Delay time indicator function for four different values of $z$ and for $E_{J}=-0.043$
@@ 6 Conclusions and final remarks
The system analysed in the present article is an instructive example of Hamiltonian dynamics with 3 not symmetry-related NHIMs, and for the effects of coordination of the bifurcation scenario of the NHIMs. A new feature in the present article is the inclusion of transient behaviour close to NHIMs generated by their break into the formation of the global scenario. We saw indications that also the transient chaotic trajectories around the NHIMs give important contributions to the global dynamics.
@@
In the study of the bifurcation scenario of the NHIMs and the comparison with the corresponding left-right symmetric system investigated in [23], we made the following interesting observation: Also, when the symmetry between the two outer NHIMs is broken by the transition to different masses of the two dwarf galaxies, there remains a partial qualitative symmetry in the bifurcation scenarios of the two outer NHIMs. This partial symmetry includes the sequence of the important bifurcations and their order for the Lyapunov orbits contained in the individual NHIMs. But it does not include the coordination of the creation of chaos. Here, in this example, only NHIM $\mathcal{M}^{1}_{E_{J}}$ and NHIM $\mathcal{M}^{3}_{E_{J}}$ make coordinated bifurcations into the direction of chaos; NHIM $\mathcal{M}^{2}_{E_{J}}$ does not participate in this coordination. This happens even though there are heteroclinic trajectories between any pair of NHIMs. The horizontal Poincaré map presented in subsection 4.2 gives at least some indication of the relatively weak heteroclinic coupling of NHIM $\mathcal{M}^{2}_{E_{J}}$ to the other two fundamental NHIMs of the system.
@@
It seems that NHIM $\mathcal{M}^{1}_{E_{J}}$ and NHIM $\mathcal{M}^{3}_{E_{J}}$ become susceptible to tangential instabilities at a similar Jacobi constant interval, while NHIM $\mathcal{M}^{2}_{E_{J}}$ is not yet ready for the development of large-scale chaos. Of course, in the symmetric case studied in [23], NHIM $\mathcal{M}^{2}_{E_{J}}$ has exactly the same bifurcation scenario as NHIM $\mathcal{M}^{3}_{E_{J}}$, and the two NHIMs develop large-scale chaos in the same Jacobi constant interval. The horizontal horseshoe plot of Fig.11 suggests that the large-scale KAM island associated with the potential minimum $P_{2}$ prevents a rapid and efficient heteroclinic connection between NHIM $\mathcal{M}^{1}_{E_{J}}$ and NHIM $\mathcal{M}^{2}_{E_{J}}$ along horizontal trajectories. In our collection of heteroclinic orbits, we do not have a single one which starts or ends on NHIM $\mathcal{M}^{2}_{E_{J}}$, runs through the potential minimum at $P_{2}$ and does not go to $z$ values far away from $z=0$. As we have seen in subsection 4.1, long and complicated trajectories can do it. As Figs.8(e) and 9(a) show, there are also trajectories going to the outside and circling around $L_{4}$ or $L_{5}$ can do it. But these long trajectories seem to be little efficient in the creation of coordination effects in the bifurcation scenarios between NHIMs. Of course, also trajectories using the vertical degree of freedom can establish heteroclinic connections, remember the trajectory from Fig.9(d). However, we have seen here (and also in other model systems) that in systems with a $z$ reflection symmetry and accordingly an invariant horizontal subsystem, the energy exchange between the vertical degree of freedom and the horizontal degrees of freedom is rather slow. This prevents an efficient heteroclinic coupling between the NHIMs via trajectories having a large part of their Jacobi constant in the vertical motion.
@@
One important aspect of the bifurcation scenario, which we can not yet explain precisely, but which certainly deserves a further, more fundamental investigation, is the following consideration: For stable and unstable manifolds of NHIMs, there holds a foliation theorem which loosely speaking says that these manifolds are the Cartesian product of a line with the internal NHIM structure. Thereby, these manifolds transport the internal NHIM structure along with them to all regions of the phase space which are heavily influenced by the homoclinic/heteroclinic tangle of the NHIMs. As a consequence, the whole homoclinic/heteroclinic tangle is at least qualitatively similar to a Cartesian product of a lower-dimensional homoclinic/heteroclinic tangle (in our case, the horizontal horseshoe presented in subsection 4.2) and the internal NHIM structures. This property has been noticed before and investigated in more detail in [36]. Also, the system investigated in this reference is 3-dof and has, in a natural form, an invariant horizontal 2-dof subsystem. Such a tendency of higher-dimensional homoclinic/heteroclinic tangles to show an approximate product structure has also been observed by another group in [37, 38, 39]. This tendency seems to be common and needs a more profound explanation.
@@
The appearance of the tangential transient effects shows that, also after the beginning of the break NHIMs, its remnants, the transient region formed close to the NHIM, have an effect on their environment in the phase space and on the formation of trajectories with similar behaviour to the homoclinic/heteroclinic tangle that existed before the break of the constituent KAM tori of the NHIMs. The fragments and transient parts continue without heavy qualitative differences.
@@
The NHIMs associated with the Lagrange points of the effective potential direct the transport of test particles through these effective potential saddles. In the interpretation of our model as describing two dwarf galaxies in interaction, this model describes the exchange of stars between the two galaxies (motion through the potential saddle associated with the central Lagrange point $L_{1}$) or the loss of stars from the binary galaxy system (motion through the potential saddles associated with the outer Lagrange points $L_{2}$ and $L_{3}$). The projection of the unstable manifolds of the NHIMs into the position space then gives the position of tidal arms and tidal bridges going out from the dwarf galaxies. Of course, our model system could also be interpreted as a simple model for two nearby globular clusters in interaction. Here, it is interesting to mention that in the astronomical literature, there is a claim of observational indications for a tidal bridge between the interacting globular clusters NGC5024 and NGC5053 [40].
Finally, let us remark that the system studied is a variant of the Lagrange three-body problem. Therefore, it has similar properties to those of other variants of the restricted three-body system. Some main qualitative properties of the phase space found in the present work are rather robust against perturbations of the system. This can be seen first from the similarity with the symmetric case studied in detail in [23]. And second, the persistence theorems for NHIMs and their invariant stable and unstable manifolds guarantee that their chaotic invariant set will be similar for a general small perturbation of the system. Small perturbations of the system might be small modifications of the gravitational potential, for example, to come to the anisotropic Kepler problem, and they could also include relativistic corrections. Also, we expect that our system has similarities with the galactic model recently studied in [41].